Project

General

Profile

Actions

HallMeeting-2014-02-05
1. Muon Beamline – Henry Nebrensky
Decay solenoid:
Pumps powered back on, going OK
Next Tuesday now for RGA clean-up; cooldown delayed by one week Andy – report delay back to Steve G Bottle pack positioned, Vicky/Mike to connect as required?
Conventional magnets: Swap over drip trays from Q7-9 to Q5-6, low priority - Mark
Henry – chase up handle on Q8 PSU
Target: Send irradiated fibres to Sheffield – Andy - standsStarted R78 target, but some variability observed, possibly investigate next Monday
Feb 3rd target run cancelled, discuss rescheduling in MEMO Request to put R78 target onto MICEnet, but be wary of security – Pierrick investigate other solutions
Proton absorber: Talk to Steve Watson about installation plan - Henry
2. FC, liquid hydrogen and absorber integration - Steve Watson
Absorber trial installation went fine, tooling worked well – indium seals and pipe routing need understanding Check mods to instrumentation cables/connections – Steve/Ian
3. Tracker activity in MICE Hall and R1/R9 – Craig Macwaters Tracker model expected next Wednesday, arrange survey through Tim Hayler – Geoff ACP quoting for spare external waveguides – Geoff Need light for darkroom – Andy G, add to list
Might need to remove darkroom base for survey
He window has been removed
4. Installation of Step IV – Andy Nichols.
West mezzanine still on target for mid-Feb completion Jason – chase start of South ramp cutting South mezzanine mod drawings late – Jason chase OPUS directly
PRY floor interface drg. complete and circulated Some minor changes to PRY design to be reported to Steve Plate – Jason Jason – request trench cable repositioning through ISIS Footprints, also find spec of ISIS control cables and susceptibility to magnet quench – Field Mitigation Group – important action! Fit brass interlayer pads to downstream SS support frame – Mark S – low priority
5. Hall Infrastructure – Andy Nichols
Lock for neutron monitor has arrived
Neutron monitor failed to power up after power outage. Frances Burge will deliver duplicate, Andys G&N investigate repairs with Lab-Impex RR2 cable channels need widening – BPG note after meeting: scheduled for next Wednesday Air conditioning exercised OK – John & Pierrick re-visit need for automation Plug beamstopper power unit back in - Henry
6. Alignment and survey – Tim Hayler Investigate reason for lateral displacement of EMR – Henry Might need to re-do Hall network - Tim
Measure tracker model in situ – Tim make request
7. Electrical Infrastructure – Andy Gallagher
Green text: action complete, delete from list
Blue text: waiting for information/parts
Red text: normal work in progress
Install power feeds to RF rack positions. Cables ordered. Drawing issued. Mark up Rack positions and have floor plates cut then install supply cables. On Hold for STEP IV installation
• Re-position 208v Transformer further east in the trench. Re-position into RR2.
• Re-position D13 on North mezz extn. Waiting on Craig’s rack position changes. Install network cables & limit cable for Proton Absorber. Remove webcam from west wall when access is available from scaffold. R9 Dark Room Lights? Install air lines and limits cable for Proton Absorber. Install cable tray to RR2. Install distribution board in RR2 when access is available. Install cable tray in RR2 when steel plinths are in position. EMR rack – install filtered air inlet and reverse fans. Install temporary supplies for PPS & Network (MICE Rack Room). Install supply cables from MICE Sub-station to RR2. This work will be done by BPG and terminated during power shutdown
8. MLCR, controls and networking, Online – Craig Macwaters/Henry Nebrensky DAQ crate’s noisy fan has failed, Craig to replace – stands More keys required for MLCR – Andy – stands
MLCR A/C supply confirmed not on substation 25 supply – chase with Gordon Stanford - Craig Change A/C filters before end FY - John
VME crates – hoovered out, two working, one failed, Craig getting quote for new equipment
miceioc6 crate has replaced miceioc1 miceonrec03 did not come back after power outage – ship to Sheffield – MOM?
Network switch UPS tested OK
9. PPS – Craig Macwaters
All OK
10. RF progress at DL and RAL – Tim Stanley
Start dismantling PSUs on March 10th for shipment to DL
Cover up equipment in Hall where practicable - Tim
11. Detector preparations and running
Nothing to report
12. R9 activity – Steve Watson One Rootes blower being tested, one remaining – Mark T
Cooldown of FC#1 going well 13K/day, at 150K now – internal leak not getting any worse
Final trial fitting of pipework tomorrow pm – Wing Lau attending
13. Effective dates of next shutdown and job list –All.
4th April – 6th May 2014, (maintenance days: 26th February, 19th March
Synchrotron phone link Neutron monitor calibration Proton absorber Drip trays for quads/dipole Clean up DSA Target chiller – talk to Paul Hodgson
14. A.O.B Web-whiteboard – needs to be updated and advertised – MOM/John G Health Physics to check dust from ESD vacuum cleaner before disposal - Henry John – investigate possible storage space at TS2
15. Date of next meeting
Wednesday, February 12th, G-06, R66, 13:30 local time.

Updated by Govans, John almost 10 years ago · 1 revisions